OpenStudy (anonymous):

please show me how to simplify this 3/8(1+1)^-5/2

OpenStudy (dumbcow):

what is being raised to the power?

OpenStudy (anonymous):

-5/2

OpenStudy (dumbcow):

no. what is being raised to the -5/2?

OpenStudy (anonymous):

nothing

OpenStudy (anonymous):

this what it is 3/8(2)^-5/2

OpenStudy (dumbcow):

is the problem written like this: (3/8)*2^-5/2 or is it written like this: ((3/8)*2)^-5/2

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

its suppose to come out to be 3/

OpenStudy (anonymous):

\[3/16\sqrt{2}\]

OpenStudy (anonymous):

can u help me on my cola problem please

OpenStudy (dumbcow):

the answer is \[\frac{3}{32\sqrt{2}}\]

OpenStudy (anonymous):

thats what i got to i dont know why the answer isnt coming out right